OpenStudy (mathisfun13):

find r for the geometric series S5=244 a1=4 and a5=324

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

a5=a1*r^4. r^4=a5/a1=324/4=81 r=+3 or r =-3 but S5<a5. a1=4, a2=4*(-3)=-12; a=3=(-12)*(-3)=+36, a4=36*(-3)=-108 a5=(-108)*(-3)=+324 a1+a2+a3+a4+a5=4-12+36-108+324=244=s5

OpenStudy (anonymous):

Thanks!

OpenStudy (mathisfun13):

so its plus or minus 3?

OpenStudy (anonymous):

Either or

OpenStudy (kropot72):

Only one of the alternatives is correct. To find which one is correct, (3 or -3), substitute each into the equation for the sum of n terms of a series: \[S _{n}=\frac{a(r ^{n}-1)}{r-1}\] for the sum of 5 terms. Only one of the alternatives gives the correct result of 244.